OpenStudy (anonymous):

Twenty of the houses on a street receive cable t.v. and seven do not. A researcher studying the amount of time that people watch t.v. intends to visit five of the houses that receive cable t.v and two that do not. How many such selections of houses are there?

OpenStudy (anonymous):

\[\dbinom{20}{5}\] \[\frac{20\times 19\times 18\times 17\times 16}{5\times 4\times 3\times 2}\]

OpenStudy (anonymous):

oh no sorry

OpenStudy (anonymous):

i should read more carefully!

OpenStudy (anonymous):

that is how many ways to pick 5 out of the 20 that receive cable. then you have to multiply by \[\dbinom{7}{2}=\frac{7\times 6}{2}=21\]

OpenStudy (anonymous):

so "final answer" is \[\dbinom{20}{5}\times \dbinom{7}{2}=15504\times 21=325584\]

OpenStudy (anonymous):

r u sure?

OpenStudy (amistre64):

would it be: 27! ----- ?? 7! 20!

OpenStudy (anonymous):

idk?

OpenStudy (anonymous):

someone else got 20

OpenStudy (amistre64):

i dont know either .... at least not 100% :) if I read it right its asking similiar to how many ways can we arrange the letters mississippi to?

OpenStudy (anonymous):

well 20 is surely wrong!

OpenStudy (anonymous):

@sensei that is a totally different kind of question. ok not "totally different" but certainly different. you are asking for arrangements

OpenStudy (amistre64):

n! --------------- maybe n1! n2! n3! ... nk!

OpenStudy (anonymous):

i am fairly certain of my answer (not like usual)

OpenStudy (amistre64):

oh im sure youre right on it ;)

OpenStudy (anonymous):

we think its to big of a selection

OpenStudy (anonymous):

counting principle says if there are n ways to do one thing and m ways to do another, then there are \[m\times n\] ways to do them both. here there are clearly \[\dbinom{20}{5}\] ways to choose 5 houses from a set of 20

OpenStudy (amistre64):

would my attempt be part of it; im thinking like the "total" part of the probability

OpenStudy (anonymous):

that is the definition of \[\dbinom{20}{5}\] and likewise there are \[\dbinom{7}{2}\] ways of selecting 2 from a set of 7

OpenStudy (amistre64):

then 5/total + 5/total .... cant seem to focus on it that well

OpenStudy (anonymous):

this is not a probability question. they are not asking for "what is the probability that mr. jones house is selected". the question is just "how many"

OpenStudy (amistre64):

ack!! .... yep

OpenStudy (anonymous):

soooo whats the answer?